OpenStudy (anonymous):

If it takes an airplane 3 hours to fly 360 miles WITH the wind and 4 hours to make the return trip AGAINST the wind, find the speed of the wind and the air speed of the plane.

OpenStudy (anonymous):

I think i know what to do...we'll see. time=distance/speed speed of plane in still air=x speed of wind=y make two equations 3=360/(x+y) 4=360(x-y) we can make the equations look a little better 3(x+y)=360 4(x-y)=360 ---------- 3x+3y=360 4x-4y=360 Now it is a system of equations where you can solve. Can you?

OpenStudy (anonymous):

Yes, now this finally makes sense! I'm going to do the equation now. ps, I apologize if it takes me a while, I broke my hand and its hard to type & write

OpenStudy (anonymous):

ill multiply the top by 4 and the bottom by 3 so I can cancel out the y

OpenStudy (anonymous):

24x = 2520

OpenStudy (anonymous):

x = 105. What does the x variable stand for though?

OpenStudy (anonymous):

wait I got it, give me a sec

OpenStudy (anonymous):

X is speed of plane in still air

OpenStudy (anonymous):

plugging in now

OpenStudy (anonymous):

y = 15! so y is the speed of the wind. thank you so much now I understand how to set it up!

OpenStudy (anonymous):

You're welcome